__________.x^3________x^2.____________x.__________________________
If x, x^2 , and x^3 lie on a number line in the order shown above, which of the following could be the value of x?
A. -2 B -1/2 C 3/4 D 1 E 3/2

Answer by CubeyThePenguin(3113)

Plug in all the numbers:

A. x^3 = -8, x^2 = 4, x = -2
B. x^3 = -1/8, x^2 = 1/4, x = -1/2
C. x^3 = 27/64, x^2 = 9/16, x = 3/4
D. x^3 = 1, x^2 = 1, x = 1
E. x^3 = 27/8, x^2 = 9/4, x = 3/2

The answer is C.